# Lost in Tracking Translation: A Comprehensive Analysis of Visual SLAM in Human-Centered XR and IoT Ecosystems

### v1

@@ Abstract.
Advancements in tracking algorithms have empowered nascent applications across various domains, from steering autonomous vehicles to guiding robots to enhancing augmented reality experiences for users. However, these algorithms are application-specific and do not work across applications with different types of motion; even a tracking algorithm designed for a given application does not work in scenarios deviating from highly standard conditions. For example, a tracking algorithm designed for robot navigation inside a building will not work for tracking the same robot in an outdoor environment. To demonstrate this problem, we evaluate the performance of the state-of-the-art tracking methods across various applications and scenarios. To inform our analysis, we first categorize algorithmic, environmental, and locomotion-related challenges faced by tracking algorithms. We quantitatively evaluate the performance using multiple tracking algorithms and representative datasets for a wide range of Internet of Things (IoT) and Extended Reality (XR) applications, including autonomous vehicles, drones, and humans. Our analysis shows that no tracking algorithm works across different applications and scenarios within applications. Ultimately, using the insights generated from our analysis, we discuss multiple approaches to improving the tracking performance using input data characterization, leveraging intermediate information, and output evaluation.
@@ 1. Introduction
Tracking systems are fundamental to immersive Extended Reality (XR) applications, facilitating accurate and real-time navigation and mapping that are crucial for creating immersive and interactive experiences (Billinghurst et al., 2015; Klein and Murray, 2007). However, various challenges must be addressed for accurate tracking, particularly in human-centered scenarios like XR (Cadena et al., 2016; Nwankwo and Rueckert, 2023). Tracking systems in XR face additional challenges due to human factors such as unpredictable movements, inter-individual variability, contextual factors, cognitive load, occlusions from body parts, physical safety concerns, adaptive requirements, and the need for real-time interaction. These elements introduce layers of unpredictability and complexity, further complicating the tracking process. These challenges are intertwined, involving (1) the environment’s complexity, (2) various locomotion demands, and (3) the inherent limitations of sensing and tracking systems. As a result, while tracking methods are often presented as generic, their performance significantly varies across different environments, locomotion scenarios, and application settings, such as drones (Burri et al., 2016), autonomous vehicles (Geiger et al., 2013), robotics (Zhang et al., 2022), and other human-centered (Ngo et al., 2022; Vávra et al., 2017; Fu et al., 2021) and non-human-centered environments (Aqel et al., 2016; Milz et al., 2018).
@@
To understand these challenges, it is essential to examine the specific factors contributing to the complexity of broader tracking systems and how human factors add to their complexity. First, the complexity of the environment can vary with the number of objects, lighting conditions, occlusions, weather, reflective surfaces, and scene changes. For instance, tracking in a crowded urban setting with changing lighting and reflective surfaces is particularly difficult. In XR applications, this complexity is heightened by the unpredictability of human interactions and the dynamic nature of the environment. Additionally, humans can seamlessly transition between different environments, such as walking from a room to a corridor to the outdoors, without a break. This continuous movement across varied settings introduces additional challenges for tracking systems, as they must constantly adapt to new conditions and maintain accuracy. Second, locomotion differs across applications. Vehicles, robots, and humans move differently, each posing unique challenges. In human-centric applications, such as XR, abrupt movements can cause blurred images. Even in-vehicle navigation typically involves fewer abrupt movements, maintaining tracking accuracy is difficult due to varying speeds and accelerations in parking lots, urban areas, and highways (Milz et al., 2018). Tracking human movement adds another layer of complexity, as humans frequently stop, walk, and change speeds unpredictably, making consistent speed maintenance challenging. Third, sensors such as IMU sensors, depth cameras, and RGB cameras each have specific issues (Chong et al., 2015). IMU sensors can drift over time (Anwar et al., 2019; Trippel et al., 2017), depth cameras struggle with lighting and reflective surfaces (Halmetschlager-Funek et al., 2018), and RGB cameras are affected by lighting variations. In XR applications, these sensor limitations are compounded by the need to integrate data from multiple sources in real-time (Ungureanu et al., 2020).
@@
To overcome these challenges, prior work has developed tracking algorithms that leverage various computational approaches. For example, traditional SLAM methods heavily depend on carefully engineered features and manually designed system components (Campos et al., 2021; Zubizarreta et al., 2020; Schops et al., 2019; Qin et al., 2018). These methods often lack robustness, meaning they struggle to maintain accuracy and reliability in dynamic and diverse real-world scenarios where conditions can vary significantly. Factors such as changing lighting conditions, moving objects, and varying environmental textures can degrade their performance. Conversely, end-to-end learning approaches (Zhou et al., 2018; Bloesch et al., 2018; Wang et al., 2020) learn system components directly from data, which can lead to improved adaptability. However, these approaches can also face robustness issues, as they may fail to generalize when encountering unfamiliar situations or environments not represented in their training data. Hybrid approaches (Tang et al., 2020; Krishna Murthy et al., 2020; Zhou et al., 2017) aim to enhance overall performance by combining traditional and learning-based methods, leveraging both strengths. While this improves the average case performance, it often sacrifices the best-case performance the individual approaches might achieve.
@@
To comprehensively address these challenges, it is important not only to evaluate tracking systems within XR environments but also to compare their performance against other application domains, such as autonomous vehicles and drones. Evaluating tracking methods across these varied domains provides a broader perspective on the strengths and weaknesses of different approaches. Autonomous vehicles and drones present unique challenges, such as high-speed movement and indoor-outdoor environmental variability, which can inform improvements in XR tracking systems. By understanding how these systems perform in different contexts, we can derive insights that contribute to developing more robust and versatile tracking solutions that can be applied across multiple domains, including but not limited to XR. Additionally, it is crucial to examine how these algorithms behaved in their original use cases (Valente et al., 2019; Zhang et al., 2021a; Milz et al., 2018; Wang et al., 2022; Chen et al., 2020) before XR became prominent. Understanding their foundational performance and limitations in traditional applications will provide a deeper insight into their adaptability and potential enhancements needed for XR environments.
@@
This paper aims to address these challenges by systematically understanding the challenges, technical requirements bottlenecks, and potential solution directions needed to enhance tracking performance in XR and beyond. In doing so, we make the following contributions:
@@
    Taxonomy of challenges. We categorize the algorithmic, environmental, and locomotion-related challenges tracking systems face and their impact on XR applications. This taxonomy provides a structured overview of the difficulties inherent in visual SLAM tracking by highlighting the specific issues that need to be addressed to improve tracking performance in various human-in-loop and other Internet of Things (IoT) systems.
@@
    Dataset characterization. Building on observations from our quantitative evaluation across traditional, end-to-end learning-based, and hybrid tracking systems, we conduct a preliminary proof of concept data characterization. This analysis highlights the importance of understanding how dataset properties impact tracking performance and identifies potential adaptive solutions for specific environments and use cases.
@@ 2. Background and Motivation
XR racking systems differ from other cyber-physical systems like autonomous vehicles, drones, and robotics due to their need to integrate virtual and real-world elements in real-time (Speicher et al., 2019). Each of these systems presents distinct tracking challenges, but XR systems face additional complexities related to human interaction (Bhargava et al., 2018; Schwind et al., 2019; Biocca et al., 2001) and environmental variability (Fua and Lepetit, 2007). Autonomous vehicles rely heavily on sensors like GPS, LiDAR, and cameras to navigate and avoid obstacles in dynamic environments (Milz et al., 2018). While real-time data processing is crucial in both XR and vehicle systems, XR tracking demands more precision and low latency to maintain user immersion and visual coherence between virtual and real-world elements (Slater et al., 1996; Wienrich et al., 2021). Unlike vehicle systems, which only focus on navigating roads and traffic (Milz et al., 2018), XR systems must blend virtual objects with the real world in a visually coherent manner, necessitating accurate spatial understanding (Wienrich et al., 2021; Sluganovic et al., 2020; Li and Mourikis, 2013; Hu et al., 2024) and low latency to maintain immersion (Slater et al., 1996).
@@
Similarly, drone navigation involves flight control, stabilization, and obstacle avoidance, often in outdoor environments with varying weather conditions and terrains (Dressel and Kochenderfer, 2019). Drones utilize GPS, IMUs, and cameras to perform tasks like localization and mapping. The primary challenges in drone tracking include maintaining stability, managing energy consumption, and ensuring safety in uncontrolled airspaces (Son et al., 2015). In contrast, XR tracking systems must handle more complex human interactions and diverse environments (Muravyev and Yakovlev, 2022; An et al., 2010; Guo et al., 2022). For example, XR applications often require users to transition seamlessly between different settings (Hübner et al., 2020), such as moving from a room to a corridor to an outdoor space (Lok, 2004; Nwankwo and Rueckert, 2023), necessitating quick adaptation to varying lighting conditions (Erickson et al., 2020), occlusions (Petit et al., 2015), and reflective surfaces (Yan et al., 2022). Robotic tracking systems, used in applications ranging from industrial automation (Saeedi et al., 2016) to service robots in healthcare (Zhang et al., 2022) and hospitality (Zhang et al., 2021b), rely on sensors like LiDAR, cameras, and ultrasonic sensors to navigate and interact with their environment. While robotic and XR tracking share similarities in sensor usage, the key difference lies in the nature of interaction. XR tracking demand higher precision and real-time processing (Lebeck et al., 2017). Additionally, XR systems must account for the unpredictability of human behavior (Hu et al., 2024), such as sudden movements and changes in speed (Nwankwo and Rueckert, 2023; Warin and Reinhardt, 2022), adding complexity not typically encountered in robotic tracking (Skarbez et al., 2020).
@@ 2.2. Unique Challenges in XR Tracking
Environmental Variability: XR systems operate in diverse environments—ranging from well-lit indoor spaces to outdoor areas with changing weather conditions. This variability demands constant recalibration without disrupting user interaction (Hübner et al., 2020; Yu et al., 2021a). XR applications often require real-time adaptation to changes in the environment or user activity, placing demands on system responsiveness (Yarossi et al., 2021).
@@ 2.3. Contributions Beyond Related Work
Previous comparative analyses of SLAM tracking methods often focus on specific methodologies or application domains, typically limited to one application such as robotics, drones, or autonomous vehicles. For instance, (Jeong et al., 2021; Wang et al., 2022) primarily explore visual odometry methods that utilize end-to-end learning, but within narrow, domain-specific contexts. In contrast, our analysis takes a more holistic approach, focusing on human-centered tracking in XR environments and highlighting the fundamental differences between conventional tracking algorithms for IoT/CPS systems and XR-based applications. By evaluating performance across three diverse datasets, two representing IoT/CPS use cases (drones, cars) and one representing an XR use case, we provide a broader perspective on the unique challenges faced in human-in-the-loop tracking systems and other domains. This comparative approach reveals distinct subtleties across different application domains, offering insights that previous studies and surveys have not comprehensively addressed, as summarized in Table 1.
First, we construct a taxonomy designed to classify challenges systematically with a exploration of algorithmic, environmental, and locomotion challenges in SLAM (§3). Unlike traditional SLAM comparisons, we explore how these factors interact within dynamic, unpredictable human-centered environments, such as those encountered in XR. We emphasize the interaction between environment and movement and its impact on methods. We bridge the gap between traditional SLAM and contemporary end-to-end learning techniques, offering a unified perspective on their evolution and interrelation. Our approach combines both qualitative insights and quantitative evaluations, leading to practical recommendations and future research directions for XR-based tracking systems.
@@ 3. Visual SLAM: Charting the Uncharted Territories
In this section, we first present an overview of the state-of-the-art tracking approaches that either leverage traditional visual Simultaneous Localization and Mapping (SLAM) methods, end-to-end learning approaches, or their combination. We also detail our key contribution of conceptualizing and devising a taxonomy of challenges. The goal of analyzing the tracking methods and categorizing their challenges is to guide our evaluation of tracking approaches in Section 4.
@@ 3.1. Visual SLAM Tracking
    A SLAM pipeline starts with data collected from sensors that observe the object and the environment. There are two types of sensors employed in SLAM applications. One type consists of proprioceptive sensors that observe a phenomenon produced and perceived within a subject. Examples of such sensors include the Inertial Measurement Unit (IMU). The second type is exteroceptive sensors, which observe stimuli external to the subject. The example of such sensors include cameras, LIDAR, and depth sensors (Chong et al., 2015).
@@
    The initial pose and map, alongside the preprocessed data, are passed to the tracking module, which is the most important stage of the SLAM pipeline. The first step of the tracking stage leverages visual odometry to detect features or landmarks in the environment. These features can be key points, descriptors, or other significant patterns. Visual odometry also tracks the features/landmarks over time and uses these features to estimate motion, i.e., pose over time (Campos et al., 2021; Qin et al., 2018; Zubizarreta et al., 2020). Keyframe-based methods identify and use key points or distinct points in the image for its operation, which can be used for matching, tracking, or reconstruction. The tracking stage is also responsible for state estimation correction, which is performed by the local bundle adjustment module by minimizing the difference between the observed position of points and their estimated position.

### v2

@@ Abstract.
Advances in tracking algorithms have enabled applications from steering autonomous vehicles to guiding robots to enhancing augmented reality experiences for users. However, these algorithms are application-specific and do not work across applications with different types of motion; even a tracking algorithm designed for a given application does not work in scenarios deviating from highly standard conditions. For example, a tracking algorithm designed for indoor robot navigation will not work for tracking the same robot in an outdoor environment. To demonstrate this problem, we evaluate the performance of the state-of-the-art tracking methods across various applications and scenarios. To inform our analysis, we first categorize environmental, locomotion-related, and algorithmic challenges faced by tracking algorithms. We quantitatively evaluate the performance using multiple tracking algorithms and representative datasets for a wide range of Internet of Things (IoT) and Extended Reality (XR) applications, including autonomous vehicles, drones, and humans. Our analysis shows that no tracking algorithm works across all applications and even in diverse scenarios within the same application. Ultimately, using the insights generated from our analysis, we discuss multiple approaches to improving the tracking performance using input data characterization, leveraging intermediate information, and output evaluation.
@@ 1. Introduction
Tracking systems are fundamental to immersive Extended Reality (XR) applications, facilitating accurate and real-time navigation and mapping that are crucial for creating immersive and interactive experiences (Billinghurst et al., 2015; Klein and Murray, 2007). However, various challenges must be addressed for accurate tracking, particularly in human-centered scenarios like XR (Cadena et al., 2016; Nwankwo and Rueckert, 2023). Tracking systems in XR face additional challenges due to human factors such as unpredictable movements, inter-individual variability, contextual factors, cognitive load, occlusions from body parts, physical safety concerns, adaptive requirements, and the need for real-time interaction. These elements introduce layers of unpredictability and complexity, further complicating the tracking process. These challenges are intertwined, involving (1) the environment’s complexity and the inherent limitations of sensing systems, (2) various locomotion demands, and (3) the inherent limitations of tracking systems. As a result, while tracking methods are often presented as generic, their performance significantly varies across different environments, locomotion scenarios, and application settings, such as drones (Burri et al., 2016), autonomous vehicles (Geiger et al., 2013), robotics (Zhang et al., 2022), and other human-centered (Ngo et al., 2022; Vávra et al., 2017; Fu et al., 2021) and non-human-centered environments (Aqel et al., 2016; Milz et al., 2018).
@@
To understand these challenges, it is essential to examine the specific factors contributing to the complexity of broader tracking systems and how human factors add to their complexity. First, the complexity of the environment can vary with the number of objects, lighting conditions, occlusions, weather, reflective surfaces, and scene changes. For instance, tracking in a crowded urban setting with changing lighting and reflective surfaces is particularly difficult, especially given varying sensing capabilities across different hardware platforms. In XR applications, this complexity is heightened by the unpredictability of human interactions and the dynamic nature of the environment. Additionally, humans can seamlessly transition between different environments, such as walking from a room to a corridor to the outdoors, without a break. This continuous movement across varied settings introduces additional challenges for tracking systems, as they must constantly adapt to new conditions and maintain accuracy. Second, locomotion differs across applications. Vehicles, robots, and humans move differently, each posing unique challenges. In human-centric applications, such as XR, abrupt movements can cause blurred images. Even in-vehicle navigation typically involves fewer abrupt movements, maintaining tracking accuracy is difficult due to varying speeds and accelerations in parking lots, urban areas, and highways (Milz et al., 2018). Tracking human movement adds another layer of complexity, as humans frequently stop, walk, and change speeds unpredictably, making consistent speed maintenance challenging. Third, sensors such as inertial measurement units (IMU) sensors, depth cameras, and RGB cameras each have specific issues (Chong et al., 2015). IMU sensors can drift over time (Anwar et al., 2019; Trippel et al., 2017), depth cameras struggle with lighting and reflective surfaces (Halmetschlager-Funek et al., 2018), and RGB cameras are affected by lighting variations. In XR applications, these sensor limitations are compounded by the need to integrate data from multiple sources in real-time (Ungureanu et al., 2020).
@@
To overcome these challenges, prior work has developed tracking algorithms that leverage various computational approaches. For example, traditional SLAM methods heavily depend on carefully engineered features and manually designed system components (Campos et al., 2021; Zubizarreta et al., 2020; Schops et al., 2019; Qin et al., 2018). These methods often lack robustness and struggle to maintain accuracy and reliability in dynamic and diverse real-world scenarios where conditions can vary significantly. Factors such as changing lighting conditions, moving objects, and varying environmental textures can degrade their performance. Conversely, end-to-end learning approaches (Zhou et al., 2018; Bloesch et al., 2018; Wang et al., 2020) learn system components directly from data, which can lead to improved adaptability. However, these approaches can also face robustness issues, as they may fail to generalize when encountering unfamiliar situations or environments not represented in their training data. Hybrid approaches (Tang et al., 2020; Krishna Murthy et al., 2020; Zhou et al., 2017) aim to enhance overall performance by combining traditional and learning-based methods, leveraging both strengths. While this improves average performance, it can reduce the peak performance of individual approaches.
@@
To comprehensively address these challenges, it is important not only to evaluate tracking systems within XR environments but also to compare their performance against other application domains, such as autonomous vehicles and drones. Evaluating tracking methods across these varied domains provides a broader perspective on the strengths and weaknesses of different approaches. Autonomous vehicles and drones present unique challenges, such as high-speed movement and indoor-outdoor environmental variability, which can inform improvements in XR tracking systems. By understanding how these systems perform in different contexts, we can derive insights that contribute to developing more robust and versatile tracking solutions that can be applied across multiple domains, including but not limited to XR. Additionally, it is crucial to examine how these algorithms behaved in their original use cases (Valente et al., 2019; Zhang et al., 2021a; Milz et al., 2018; Wang et al., 2022a; Chen et al., 2020) before XR became prominent. Understanding their foundational performance and limitations in traditional applications will provide a deeper insight into their adaptability and potential enhancements needed for XR environments.
@@
This paper aims to address these challenges by systematically understanding the challenges, technical requirements, bottlenecks, and potential solution directions needed to enhance tracking performance in XR and beyond. In doing so, we make the following contributions:
@@
    Taxonomy of challenges. We categorize the algorithmic, environmental, and locomotion-related challenges that tracking systems face and their impact on XR applications. This taxonomy provides a structured overview of the difficulties inherent in Visual SLAM tracking by highlighting the specific issues that need to be addressed to improve tracking performance in various human-in-loop and other Internet of Things (IoT) systems.
@@
    Dataset characterization. Building on observations from our quantitative evaluation across traditional, end-to-end learning-based, and hybrid tracking systems, we conduct a preliminary proof-of-concept data characterization. This analysis highlights the importance of understanding how dataset properties impact tracking performance and identifies potential adaptive solutions for specific environments and use cases.
@@ 2. Background and Motivation
XR tracking systems differ from other cyber-physical systems like autonomous vehicles, drones, and robotics due to their need to integrate virtual and real-world elements in real-time (Speicher et al., 2019). Each of these systems presents distinct tracking challenges, but XR systems face additional complexities related to human interaction (Bhargava et al., 2018; Schwind et al., 2019; Biocca et al., 2001) and environmental variability (Fua and Lepetit, 2007). Autonomous vehicles rely heavily on sensors like GPS, LiDAR, and cameras to navigate and avoid obstacles in dynamic environments (Milz et al., 2018). While real-time data processing is crucial in both XR and vehicle systems, XR tracking demands more precision and low latency to maintain user immersion and visual coherence between virtual and real-world elements (Slater et al., 1996; Wienrich et al., 2021). Unlike vehicle systems, which only focus on navigating roads and traffic (Milz et al., 2018), XR systems must blend virtual objects with the real world in a visually coherent manner, necessitating accurate spatial understanding (Wienrich et al., 2021; Sluganovic et al., 2020; Li and Mourikis, 2013; Hu et al., 2024) and low latency to maintain immersion (Slater et al., 1996).
@@
Similarly, drone navigation involves flight control, stabilization, and obstacle avoidance, often in outdoor environments with varying weather conditions and terrains (Dressel and Kochenderfer, 2019). Drones utilize GPS, IMUs, and cameras to perform tasks like localization and mapping. The primary challenges in drone tracking include maintaining stability, managing energy consumption, and ensuring safety in uncontrolled airspaces (Son et al., 2015). In contrast, XR tracking systems must handle more complex human interactions and diverse environments (Muravyev and Yakovlev, 2022; An et al., 2010; Guo et al., 2022). For example, XR applications often require users to transition seamlessly between different settings (Hübner et al., 2020), such as moving from a room to a corridor to an outdoor space (Lok, 2004; Nwankwo and Rueckert, 2023), necessitating quick adaptation to varying lighting conditions (Erickson et al., 2020), occlusions (Petit et al., 2015), and reflective surfaces (Yan et al., 2022). Robotic tracking systems, used in applications ranging from industrial automation (Saeedi et al., 2016) to service robots in healthcare (Zhang et al., 2022) and hospitality (Zhang et al., 2021b), rely on sensors like LiDAR, cameras, and ultrasonic sensors to navigate and interact with their environment. While robotic and XR tracking share similarities in sensor usage, the key difference lies in the nature of interaction. XR tracking demands higher precision and real-time processing (Lebeck et al., 2017). Additionally, XR systems must account for the unpredictability of human behavior (Hu et al., 2024), such as sudden movements and changes in speed (Nwankwo and Rueckert, 2023; Warin and Reinhardt, 2022), adding complexity not typically encountered in robotic tracking (Skarbez et al., 2020).
@@ 2.2. Unique Challenges in XR Tracking
Environmental Variability: XR systems operate in diverse environments, ranging from well-lit indoor spaces to outdoor areas with changing weather conditions. This variability demands constant recalibration without disrupting user interaction (Hübner et al., 2020; Yu et al., 2021a). XR applications often require real-time adaptation to changes in the environment or user activity, placing demands on system responsiveness (Yarossi et al., 2021).
@@ 2.3. Contributions Beyond Related Work
Previous comparative analyses of SLAM tracking methods often focus on specific methodologies or application domains, typically limited to one application such as robotics, drones, or autonomous vehicles. For instance, (Jeong et al., 2021; Wang et al., 2022a) primarily explore visual odometry [^1] Visual Odometry (VO) estimates the camera’s motion (trajectory) over time by analyzing sequential visual data, without explicitly building or maintaining a global map of the environment. methods that utilize end-to-end learning, but within narrow, domain-specific contexts. In contrast, our analysis takes a more holistic approach, focusing on human-centered tracking in XR environments and highlighting the fundamental differences between conventional tracking algorithms for IoT/CPS systems and XR-based applications. By evaluating performance across three diverse datasets, two representing IoT/CPS use cases (drones, cars) and one representing an XR use case, we provide a broader perspective on the unique challenges faced in human-in-the-loop tracking systems and other domains. This comparative approach reveals distinct subtleties across different application domains, offering insights that previous studies and surveys have not comprehensively addressed, as summarized in Table 1.
First, we construct a taxonomy designed to classify challenges systematically with an exploration of algorithmic, environmental, and locomotion challenges in SLAM (§3). Unlike traditional SLAM comparisons, we explore how these factors interact within dynamic, unpredictable human-centered environments, such as those encountered in XR. We emphasize the interaction between the environment and movement and its impact on methods. We bridge the gap between traditional SLAM and contemporary end-to-end learning techniques, offering a unified perspective on their evolution and interrelation. Our approach combines both qualitative insights and quantitative evaluations, leading to practical recommendations and future research directions for XR-based tracking systems.
@@ 3. Visual SLAM: Charting the Uncharted Territories
In this section, we first present an overview of the state-of-the-art tracking approaches that either leverage traditional visual Simultaneous Localization and Mapping (SLAM) methods, end-to-end learning approaches, or their combination. Visual SLAM refers to the task of estimating a camera’s trajectory (pose) while simultaneously constructing a map of the surrounding environment, using visual input such as monocular or stereo images. Unlike conventional SLAM systems that may integrate additional sensors (LiDAR, IMU), Visual SLAM relies primarily on visual data and is widely used in robotics, AR/VR/XR, and IoT applications. We also detail our key contribution of conceptualizing and devising a taxonomy of challenges. The goal of analyzing the tracking methods and categorizing their challenges is to guide our evaluation of tracking approaches in Section 4.
@@ 3.1. Visual SLAM Tracking
    A SLAM pipeline starts with data collected from sensors that observe the object and the environment. There are two types of sensors employed in SLAM applications. One type consists of proprioceptive sensors that observe a phenomenon produced and perceived within a subject. Examples of such sensors include the Inertial Measurement Unit (IMU). The second type is exteroceptive sensors, which observe stimuli external to the subject. The example of such sensors includes cameras, LIDAR, and depth sensors (Chong et al., 2015).
@@
    The initial pose and map, alongside the preprocessed data, are passed to the tracking module, which is the most important stage of the SLAM pipeline. The first step of the tracking stage leverages visual odometry to detect and track features or landmarks in the environment. These features can be keypoints, descriptors, or other significant patterns, and are used to estimate motion (pose over time (Campos et al., 2021; Qin et al., 2018; Zubizarreta et al., 2020)). Feature-based methods identify and use keypoints or distinct image features for matching, tracking, and reconstruction. In parallel, keyframe-based methods select a subset of frames as keyframes to incrementally build the map and perform bundle adjustment more efficiently. This reduces computational cost compared to using all frames, making real-time operation feasible on mobile systems (Schöps et al., 2019; Klein, 2019). The tracking stage is also responsible for state estimation correction, which is performed by the local bundle adjustment module by minimizing the difference between the observed position of points and their estimated position.